A teacher realizes that his students' marks on a statistic test are normally distributed with a mean of 62 and a standard deviation of 15. If the teacher wishes to assign minimum marks for grade B is referring to the top 30 per cent of the students' marks in the class, how many marks are required to get a minimum mark for grade B?


1
Expert's answer
2021-04-15T07:52:36-0400

"μ=62 \\

σ=15"

Determine the z-score in the normal probability that corresponds with the probability of 1 - 0.3 = 0.70:

"z= 0.52 \\

z = \frac{x-μ}{σ} \\

x = μ + zσ \\

x = 62 + 0.52 \times 15 = 69.8 ≈70"

So, 70 marks are required to get a minimum mark for grade B